# Can Tool-augmented Large Language Models be Aware of Incomplete Conditions?

### v2

@@ 5 Results
6 Analysis
We perform an analysis to understand how LLMs perceive and respond in incomplete conditions.
6.1 Human Analysis
@@
The results of the human analysis, as presented in Table 5, indicate that the average accuracy achieved by human participants is 81.11%, with an F1 score of 80.0%. Compared to these human performance metrics, most open-source models underperform, while larger models, such as GPT-4, demonstrate comparable or superior performance. This indicates a clear benchmark for model evaluation and demonstrates the competitive performance of larger models in handling incomplete situations.
@@
6.2 Implicit Evaluation Results with Free-form Generation
As described in Section 4, our experimental setup involves a binary classification task in which models are asked to determine whether the current conditions are incomplete for tool invocation by answering "Yes" or "No." This approach directly assesses the models’ ability to evaluate whether the necessary conditions for tool usage are met. We also consider an alternative method that implicitly gauges the models’ awareness to complement our explicit evaluation. In this setup, the models engage in a simulated dialogue with a user, and their next actions are monitored. If the model generates a response that includes invoking a tool, it is considered the model has determined that it can invoke the tool. In contrast, the absence of tool invocation is considered that the model has determined that it cannot invoke the tool. We apply this implicit evaluation framework to GPT-3.5-Turbo and GPT-4 on the APIBank dataset. The text prompt used in the experiment is illustrated in Fig. LABEL:fig:app:free-form.
Utterance Errors in GPT-3.5-T GPT-4
Removal API Call
96.0 86.0
7.0 67.0
13.0 99.0
Table 7: The results of providing API invocation outcomes to LLMs across different scenarios. Accuracy is used as the evaluation metric, highlighting the differences in how GPT-3.5-Turbo and GPT-4 handle erroneous API call results and their ability to recognize hallucinations and incomplete conditions.
6.3 Impact of Tool Invocation Feedback on Incomplete Condition Recognition
Our experiment primarily focuses on assessing whether LLMs can recognize incomplete conditions before invoking a tool. This approach is practical as it helps mitigate potential safety or security risks associated with incorrect tool invocation. However, in more relaxed environments where some degree of incorrect tool usage is acceptable, allowing the model to observe the outcomes of tool invocation and reassess the completeness of the conditions could potentially enhance performance. Recent studies in tool-agent frameworks also suggest that observing the outcomes of previous actions and adjusting plans accordingly can improve task execution (react; qin2023toolllm). Building on these insights, we explore whether providing LLMs with the results of tool invocation can enhance their ability to recognize incomplete conditions.
@@ 6.3 Impact of Tool Invocation Feedback on Incomplete Condition Recognition
When API call results are erroneous, GPT-3.5-Turbo often concludes that the API cannot be invoked, regardless of whether the necessary conditions for a successful API call are met, as demonstrated in Table 7. In contrast, GPT-4 shows a stronger ability to assess the feasibility of API calls accurately, even in complex scenarios. Notably, when the necessary information to invoke APIs is missing, but the API call is still successful, GPT-3.5-Turbo exhibits limited awareness of hallucinations. On the other hand, GPT-4 is more adept at recognizing hallucinated API calls and correctly identifying incomplete conditions, even when the API call succeeds. These results may suggest that the reasoning ability of LLMs plays a critical role in interpreting API call results and making accurate predictions.
@@
Table 8: Performance comparison between Llama-2 model family (touvron2023llama) of different sizes and training approaches. Models not trained on instruction-tuning datasets are marked as BASE.
@@
6.4 Impacts of Model Size and Instruction Tuning

### v3

@@ 5 Results and Analysis
6 Conclusion
In this work, we investigate the capability of tool-augmented LLMs to recognize incomplete conditions. Specifically, we examine scenarios where users only provide partial information or the required tools are inaccessible. Our experimental results, based on carefully manipulated datasets, reveal varying performance patterns across different incomplete scenarios and model scales. While larger models show promising results, the effectiveness of enhancement strategies like few-shot learning and Chain-of-Thought prompting varies across different manipulation types and datasets. These findings highlight the importance of considering diverse incomplete scenarios in developing more reliable and safer tool-augmented LLMs for real-world applications.
Limitations
While we explore the ability of tool-augmented LLMs to abstain from tool usage in incomplete conditions, our study has several limitations. First, the data annotation process primarily relies on model-based annotations, although human verification is conducted. This approach might differ from actual cases where humans provide incomplete information regarding tool usage. Additionally, our focus is solely on API-based tools, which, although significant, do not encompass the full spectrum of tools such as plugins, robotic systems, and other interactive systems. Despite these limitations, our research underscores the importance of developing reliable tool-augmented LLMs, highlighting the need for further advancements in this area.
Appendix A Experimental Details
A.1 Implementation Details
All open-source LLMs are implemented with the Transformers library (Wolf et al., 2020). For proprietary models, we use claude-3-haiku-20240307, gpt-3.5-turbo-0125, gpt-4-0613, gpt-4o-mini-2024-07-18, and gpt-4o-2024-11-20 for Claude-3-Haiku, GPT-3.5-Turbo, GPT-4, GPT-4o-mini, and GPT-4o, respectively. The temperature is set to 0 across all models. We use two original and two manipulated instances to evaluate with few-shot examples. In experiments with instances from ToolBench, we select models with a context length exceeding 8192 tokens to handle the frequently encountered lengthy API descriptions.
A.2 Detailed Dataset Verification Process
Our preliminary analysis revealed that automatic generation sometimes failed to create truly incomplete scenarios, making manual verification crucial for maintaining dataset quality. To resolve this, we conducted the dataset verification process involved rigorous criteria for both manipulation types. For Relevant API Replacement, we removed instances where:
  • The user’s request could be fulfilled using alternative APIs
  • The request could be handled using non-replaced APIs, particularly in ToolBench instances
For Partial Removal of User Utterance, we filtered out cases where:
  • Only non-essential information was removed, allowing APIs to still fulfill the request
  • One of multiple requests was removed, but remaining requests could be resolved through existing API invocations
  • The conversation context became unnatural or grammatically incorrect
  • The result was identical to the original instance
  • The last utterance no longer pertained to the user, making API execution unnatural
Two authors independently reviewed the instances, discussing and resolving any disagreements to ensure consistent application of these criteria. This process was iteratively refined until we achieved a high level of agreement on the identification of truly incomplete scenarios.
A.3 Human Analysis
@@
Appendix B Further Analysis
We perform analyses to understand how LLMs perceive and respond in incomplete conditions.
B.1 Implicit Evaluation Results with Free-form Generation
As described in Section 4, our experimental setup involves a binary classification task in which models are asked to determine whether the current conditions are incomplete for tool invocation by answering "Yes" or "No." This approach directly assesses the models’ ability to evaluate whether the necessary conditions for tool usage are met. We also consider an alternative method that implicitly gauges the models’ awareness to complement our explicit evaluation. In this setup, the models engage in a simulated dialogue with a user, and their next actions are monitored. If the model generates a response that includes invoking a tool, it is considered the model has determined that it can invoke the tool. In contrast, the absence of tool invocation is considered that the model has determined that it cannot invoke the tool. We apply this implicit evaluation framework to GPT-3.5-Turbo and GPT-4 on the APIBank dataset.
@@ B.1 Implicit Evaluation Results with Free-form Generation
Utterance Errors in GPT-3.5 GPT-4 GPT-4o GPT-4o Removal API Call Turbo mini 96.0 86.0 97.0 96.0 7.0 67.0 45.0 79.0 13.0 99.0 51.0 91.0
Table 5: The results of providing API invocation outcomes to LLMs across different scenarios. Accuracy is used as the evaluation metric, highlighting the differences in how different models handle erroneous API call results and their ability to recognize hallucinations and incomplete conditions.
B.2 Impact of Tool Invocation Feedback on Incomplete Condition Recognition
Our experiment primarily focuses on assessing whether LLMs can recognize incomplete conditions before invoking a tool. This approach is practical as it helps mitigate potential safety or security risks associated with incorrect tool invocation. However, in more relaxed environments where some degree of incorrect tool usage is acceptable, allowing the model to observe the outcomes of tool invocation and reassess the completeness of the conditions could potentially enhance performance. Recent studies in tool-agent frameworks also suggest that observing the outcomes of previous actions and adjusting plans accordingly can improve task execution (Yao et al., 2023; Qin et al., 2023). Building on these insights, we explore whether providing LLMs with the results of tool invocation can enhance their ability to recognize incomplete conditions.
@@ B.2 Impact of Tool Invocation Feedback on Incomplete Condition Recognition
When API call results are erroneous, GPT-3.5-Turbo often concludes that the API cannot be invoked, regardless of whether the necessary conditions for a successful API call are met, as demonstrated in Table 5. In contrast, GPT-4 and GPT-4o show a stronger ability to assess the feasibility of API calls accurately, even in complex scenarios. Notably, when the necessary information to invoke APIs is missing, but the API call is still successful, GPT-3.5-Turbo and GPT-4o-mini exhibit limited awareness of hallucinations. On the other hand, GPT-4 and GPT-4o is more adept at recognizing hallucinated API calls and correctly identifying incomplete conditions, even when the API call succeeds. These results may suggest that the reasoning ability of LLMs plays a critical role in interpreting API call results and making accurate predictions.
@@
Table 6: Performance comparison between Llama-2 model family (Touvron et al., 2023) of different sizes and training approaches. Models not trained on instruction-tuning datasets are marked as BASE.
@@
B.3 Impacts of Model Size and Instruction Tuning